# TrioXpert: An automated incident management framework for microservice system

### v1

@@ Abstract
Automated incident management plays a pivotal role in large-scale microservice systems. However, many existing methods rely solely on single-modal data (e.g., metrics, logs, and traces) and struggle to simultaneously address multiple downstream tasks, including anomaly detection (AD), failure triage (FT), and root cause localization (RCL). Moreover, the lack of clear reasoning evidence in current techniques often leads to insufficient interpretability. To address these limitations, we propose TrioXpert, an end-to-end incident management framework capable of fully leveraging multimodal data. TrioXpert designs three independent data processing pipelines based on the inherent characteristics of different modalities, comprehensively characterizing the operational status of microservice systems from both numerical and textual dimensions. It employs a collaborative reasoning mechanism using large language models (LLMs) to simultaneously handle multiple tasks while providing clear reasoning evidence to ensure strong interpretability. We conducted extensive evaluations on two popular microservice system datasets, and the experimental results demonstrate that TrioXpert achieves outstanding performance in AD (improving by 4.7% to 57.7%), FT (improving by 2.1% to 40.6%), and RCL (improving by 1.6% to 163.1%) tasks.
Index Terms:
Anomaly Detection, Failure Triage, Root Cause Localization, Microservice System
@@ I Introduction
Consider a real-world incident in a microservice system where a sudden drop in throughput and success rate is observed (Fig. 1). The anomaly detection stage identifies this deviation and raises an alert for the time window “[2025.02.01 19:00:00, 2025.02.01 19:05:30]”, marking the start of the incident management. Further analysis of associated logs reveals multiple “error” and “warning” level entries during the same period, leading the incident to be classified as a “Database Connection Timeout” and assigned to the database team.
The database team then leverages trace data to examine service dependencies and identifies “Dbservice-3” as the root cause instance. By increasing its connection pool limit and restarting the instance, the team successfully repairs the incident and restores system functionality.
Among the four stages, anomaly detection, failure triage, and root cause localization are both critical and amenable to automation, making them the primary targets for AI-driven solutions. In contrast, mitigation typically involves manual intervention and complex orchestration, thus remains out of scope for this work. We therefore focus on advancing automation across the first three stages.
@@
Even if we design a better framework integrated LLMs to address these limitations, intrinsic constraints of LLMs pose the third challenge to its effectiveness in real-world deployment. Internally, issues such as hallucinations and context window limits are problems that persist regardless of model prompting or fine-tuning strategies. These limitations further prevent models from generating stable, rigorous reasoning paths that satisfy the interpretability requirements of incident management. We summarize the challenges as follows, and more details are provided in Section II:
@@
To tackle the aforementioned challenges, this paper proposes TrioXpert, an end-to-end incident management framework based on LLMs collaboration. Specifically, TrioXpert consists of three core modules: (1) Multimodal Data Preprocessing: To handle the heterogeneity of metrics, logs, and traces, TrioXpert employs three modality-specific preprocessing pipelines and two filtering mechanisms to extract incident-relevant logs and traces, reducing noise and improving data quality. (2) Multi-Dimensional System Status Representation: This module consolidates numerical and textual features into a unified multi-view representation, capturing statistical patterns, semantic context, and service interactions. (3) LLMs Collaborative Reasoning: A Mixture-of-Experts (MoE) architecture coordinates three LLM-based agents—Numerical, Textual, and Incident Experts—to jointly perform AD, FT, and RCL. Outputs include structured reasoning chains to enhance interpretability and trust. The contributions of this work are summarized as follows:
@@
    We propose TrioXpert, the first end-to-end framework that unifies AD, FT, and RCL using a collaborative reasoning architecture built on LLMs. The design integrates metrics, logs, and traces through a structured Mixture-of-Experts prompting strategy, enabling scalable and explainable multi-task diagnostics.
@@
    We conducted extensive experiments on two popular microservice systems. The results demonstrate that TrioXpert outperforms baseline methods in terms of performance and interpretability. Furthermore, ablation studies were performed to further validate the effectiveness and necessity of each module design in this work. Our source code and experimental data are publicly available  [^1] https://anonymous.4open.science/r/TrioXpert-F244.
@@ II-A1 Do metrics, logs, and traces all carry diagnostic value
To understand how various modalities characterize microservice systems, we conducted a qualitative and quantitative analysis on dataset $\mathscr{D}2$. We randomly sampled 20 incident cases and collected corresponding metrics, raw logs, and traces within a 10-minute window around each incident.
@@ II-C Problem Definition
Typically, OCEs need to collect multimodal data (i.e., metrics, logs, and traces) and perform three core tasks: AD, FT, and RCL. Specifically, the goal of the AD task is to detect whether there are multiple abnormal timestamps within a time window, thereby determining whether the system has encountered an anomaly. The FT task involves selecting the most probable failure type $c_{i}$ from a predefined set of fault categories =$\mathbb{C}=\{c_{1},c_{2},...,c_{n}\}$. Finally, the RCL task identifies the most likely root cause $r_{i}$ from all service instances =$\mathbb{R}=\{r_{1},r_{2},...,r_{m}\}$.
@@ III-A Overview
In the preprocessing stage, tailored strategies are applied to metrics, logs, and traces to produce structured inputs: a time-series matrix $\mathcal{M}$, a service topology graph $\mathcal{G}$, and filtered logs $\mathcal{L}$ and traces $\mathcal{T}$. Subsequently, three modality-aware pipelines extract numerical and textual features for system status characterization. Finally, a MoE architecture coordinates multiple LLMs to jointly perform downstream incident management tasks, supported by structured reasoning evidence.
Figure 2: The overview of TrioXpert. The framework consists of three modules: (1) Multimodal Data Preprocessing, which filters and prepares metrics, traces, and logs (e.g., aligning by time and extracting relevant subsets); (2) Multi-Dimensional System Status Representation, which derives a numerical feature-based preliminary analysis and textual summaries of logs/traces (via specialized “Log Summarizer” and “Trace Summarizer”); and (3) LLMs Collaborative Reasoning, a Mixture-of-Experts comprising a Numerical Expert and Textual Expert that collaborate with an Incident Expert LLM to perform anomaly detection, failure triage, and root cause localization simultaneously.
Furthermore, the overall framework of TrioXpert comprises two key phases: offline training and online testing. During the offline training phase, the model utilizes data from normal periods to train the feature processor in the first data processing pipeline, learning the normal feature representations. In the online testing phase, the model sequentially processes data through the aforementioned three modules, generating results that are both accurate and supported by detailed reasoning.
@@ III-B1 Metrics
Considering the time-series characteristics of metrics, it is organized into a three-dimensional time-series matrix $\mathcal{M}$. The three dimensions of this matrix represent timestamps, service instances, and feature channels, respectively. This structured representation not only facilitates subsequent time-series feature extraction but also effectively captures the dynamic changes in system operational status.
@@ III-B2 Logs
In the first stage, the LLM is prompted to extract incident-relevant keywords from the dataset (e.g., generating terms such as “error”, “failure”, or “timeout” in response to the instruction “Identify key terms commonly associated with system failures in the following logs” ). These keywords are then employed to filter candidate logs, which are subsequently fed into the second stage for deeper semantic analysis. In the second stage, the LLM identifies salient log entries by analyzing their semantic content in the context of incident patterns, such as whether they contain abnormal operations, error indicators, or system-critical events. This process replaces the TF-IDF-based prioritization used in traditional approaches with a more expressive, semantics-aware mechanism.
@@
Definition 2: Filtered logs $\mathcal{L}$ are a refined subset of log entries obtained through a two-stage process: keyword-based filtering to retain incident-related logs, followed by TF-IDF prioritization to select distinctive entries.
@@ III-B3 Traces
Then, due to the varying latency distribution characteristics of different invocation types (e.g., HTTP, RPC) in trace data, we independently calculates the P95 latency percentile for all spans of each invocation type as the threshold. For spans with latency exceeding this threshold, their parent spans are recursively traced back to the root node to ensure the preservation of complete invocation chains. This approach supports contextual analysis of anomalous invocation chains while removing irrelevant trace data. Additionally, following the methods of ART[1] and DiagFusion[2], TrioXpert extracts service topology graphs $\mathcal{G}$ from trace data, maintaining the same sampling frequency as the metrics data to describe service instance invocation relationships at any given time.
@@ III-C Multi-Dimensional System Status Representation
To comprehensively capture the state of a microservice system, TrioXpert employs three independent data processing pipelines, targeting both numerical and textual modalities. Each pipeline extracts and encodes salient features from a specific modality, providing structured inputs for downstream reasoning.
@@ III-C1 Numerical Perspective
The numerical pipeline processes the time-series metric matrix $\mathcal{M}$ and the corresponding service topology graph $\mathcal{G}$. A feature processor predicts the next-step values of $\mathcal{M}$ and computes the deviation from ground truth. By minimizing the L2 norm of this deviation during training, the model learns robust representations of normal behavior in a self-supervised manner. Deviations are then used to generate preliminary answers. This pipeline draws inspiration from the ART [1] methodology, employing joint analysis of time-series data and topological relationships to provide a comprehensive description of the microservice system’s operational status from the perspective of numerical features.
@@ III-D LLMs Collaborative Reasoning
To address these challenges, TrioXpert adopts the MoE architecture comprising three specialized LLMs: a Numerical Expert, a Textual Expert, and an Incident Expert. Each expert operates on modality-specific inputs and contributes to a collaborative diagnostic process.
@@ III-D2 Coordination pipeline
TrioXpert adopts the MoE paradigm, leveraging a dispatch-and-aggregation mechanism to fully exploit the specialized strengths of each expert in their respective domains. The pseudo-code for the coordination pipeline is shown in Algorithm 1. Specifically, this mechanism first assigns numerical features $\mathcal{F}_{num}$ and textual features $\mathcal{F}_{txt}$ to the Numerical Expert and Textual Expert, respectively, for independent analysis. Both experts independently generate results and corresponding reasoning evidence for AD, FT, and RCL. Subsequently, the Incident Expert performs a comprehensive analysis of the outputs from the Numerical Expert and Textual Expert, balancing information from different sources to resolve potential conflicts and ultimately producing a unified and reliable diagnostic result $\mathcal{R}_{final}$ along with comprehensive reasoning evidence $\mathcal{E}_{final}$. This dispatch-and-aggregation mechanism not only enables efficient integration of multimodal data but also significantly enhances the accuracy and interpretability of TrioXpert in performing AD, FT, and RCL tasks within complex microservice systems.
@@ 2 $\mathcal{R}_{txt},\mathcal{E}_{txt}\leftarrow$ TextualExpert($\mathcal{F}_{txt}$) ;
Algorithm 1 LLM-Based Coordination Pipeline in TrioXpert
@@ III-D3 Conflict resolution and aggregation policy
Due to potential inconsistencies in incident information across modalities, conflicts may arise between experts. For example, the Numerical Expert may infer a failure type such as “Node CPU overload”, based on time-series metrics, while the Textual Expert may conclude it is a “Container Hardware” issue based on logs and traces. To resolve such discrepancies and generate robust decisions, we introduce a weighted fusion mechanism that prioritizes more reliable data sources for each subtask.
@@ III-D4 Hallucination mitigation design
To address the challenge of hallucination in LLMs, TrioXpert introduces a comprehensive mitigation strategy centered around refined prompt engineering. Specifically, TrioXpert adopts a structured prompt engineering approach grounded in the “RGCIE” principle (i.e., Role, Goal, Constraints, Instructions, Example) to tailor prompts for different LLMs. By explicitly defining roles, goals, instructions, and constraints, along with providing high-quality examples, this approach ensures that LLMs strictly adhere to task requirements, thereby significantly reducing the occurrence of hallucinations. As shown in Fig. 4The prompts are further formalized using YAML format, enabling consistent structuring, easy customization, and programmable parsing across diverse model interfaces. This systematic design not only enhances the precision of model outputs but also aligns them more closely with the intended objectives.
Through this strategy, TrioXpert effectively mitigates the impact of hallucination, ensuring robust and reliable performance in incident management tasks.
@@ IV Experiments
    RQ3: Does TrioXpert generates interpretable results?
@@ IV-A1 Datasets
To comprehensively evaluate the performance of TrioXpert in AD, FT, and RCL, we conducted extensive experiments on two multimodal datasets, $\mathscr{D}1$ and $\mathscr{D}2$. These datasets include three modalities of data (i.e., metrics, logs, traces), and each incident case is explicitly annotated with the timestamp of failure injection, the failure type, and the corresponding root cause service instance. Since these datasets not only encompass multimodal information but also provide high-quality annotations consistent with multi-task labels, they effectively support a comprehensive evaluation of TrioXpert’s performance.
@@ IV-A3 Evaluation Metrics
As described in Section II-C, the goal of TrioXpert is to detect failures, determine failure types, and locate the root cause. To better reflect the performance of the evaluated methods in real-world applications, we carefully designed and selected different evaluation metrics for each task.
AD and FT are both inherently classification tasks, but their objectives differ: AD is a binary classification task used to determine whether a system has experienced a failure, while FT is a multi-class classification task aimed at identifying the type to which the failure belongs. To evaluate the performance of these two tasks, we utilized the following metrics: True Positives (TP), False Positives (FP), and False Negatives (FN). Based on these metrics, we calculated $Precision$ and $Recall$, and subsequently derived the $F1$ as a comprehensive evaluation criterion.
$Precision=\frac{TP}{TP+FP}$ (1)
$Recall=\frac{TP}{TP+FN}$
$F1=\frac{2\cdot Precision\cdot Recall}{(Precision+Recall)}$
For the RCL task, we introduced the $Top@K$ metric to calculate the probability that the true root cause is included within the top $K$ predicted results, where $gt$ represents groundtruth, $P$ stands for predictions and $K$ is the total amount of failures. Additionally, to comprehensively assess model performance, we also employed the $Avg@5$ metric.
$Top@K=\frac{1}{N}\sum_{i=1}^{N}(gt_{i}\in P_{i}[1:K])$ (2)
$Avg@5=\frac{1}{5}\sum_{i=1}^{5}Top@K$
@@
IV-A4 Backbone Selection
We choose Qwen2.5-7B-Instruct as the backbone LLM in our framework, as it achieves state-of-the-art performance on the Open LLM Leaderboard [30] among models with fewer than 10 billion parameters.
IV-A5 Implementations
TrioXpert and the baseline methods were implemented using Python 3.10.16, PyTorch 2.4.0, Transformers 4.46.0, DGL 2.4.0+cu124, and Accelerate 1.4.0. The experiments were conducted on a server equipped with a 16-core Intel(R) Xeon(R) Gold 5416S CPU, 376 GB of RAM, and 8 NVIDIA RTX A6000 GPUs, each with 48 GB of memory. To minimize the impact of randomness on the experimental results, each experiment was repeated five times, and calculate the average value as the final result.
IV-B RQ1: How does TrioXpert perform in AD, FT, and RCL?

### v2

@@ Abstract
Automated incident management plays a pivotal role in large-scale microservice systems. However, many existing methods rely solely on single-modal data (e.g., metrics, logs, and traces) and struggle to simultaneously address multiple downstream tasks, including anomaly detection (AD), failure triage (FT), and root cause localization (RCL). Moreover, the lack of clear reasoning evidence in current techniques often leads to insufficient interpretability. To address these limitations, we propose TrioXpert, an end-to-end incident management framework capable of fully leveraging multimodal data. TrioXpert designs three independent data processing pipelines based on the inherent characteristics of different modalities, comprehensively characterizing the operational status of microservice systems from both numerical and textual dimensions. It employs a collaborative reasoning mechanism using large language models (LLMs) to simultaneously handle multiple tasks while providing clear reasoning evidence to ensure strong interpretability. We conducted extensive evaluations on two microservice system datasets, and the experimental results demonstrate that TrioXpert achieves outstanding performance in AD (improving by 4.7% to 57.7%), FT (improving by 2.1% to 40.6%), and RCL (improving by 1.6% to 163.1%) tasks. TrioXpert has also been deployed in Lenovo’s production environment, demonstrating substantial gains in diagnostic efficiency and accuracy.
@@ I Introduction
Among the four stages, AD, FT, and RCL are both critical and amenable to automation, making them the primary targets for AI-driven solutions. In contrast, mitigation typically involves manual intervention and complex orchestration, thus remains out of scope for this work. We therefore focus on advancing automation across the first three stages.
@@
Even if we design a better framework integrating LLMs to address these limitations, intrinsic constraints of LLMs pose the third challenge to its effectiveness in real-world deployment. Internally, issues such as hallucinations and context window limits are problems that persist regardless of model prompting or fine-tuning strategies. These limitations further prevent models from generating stable, rigorous reasoning paths that satisfy the interpretability requirements of incident management. We summarize the challenges as follows, and more details are provided in Section II:
@@
To tackle the aforementioned challenges, this paper proposes TrioXpert, an end-to-end incident management framework based on LLMs collaboration. Specifically, TrioXpert consists of three core modules: (1) Multimodal Data Preprocessing: To handle the heterogeneity of metrics, logs, and traces, TrioXpert employs three modality-specific preprocessing pipelines and two filtering mechanisms to extract incident-relevant logs and traces, reducing noise and improving data quality. (2) Multi-Dimensional System Status Representation: This module consolidates numerical and textual features into a unified multi-view representation, capturing statistical patterns, semantic context, and service interactions. (3) LLMs Collaborative Reasoning: An multi-expert collaboration architecture coordinates three LLM-based agents—Numerical, Textual, and Incident Experts—to jointly perform AD, FT, and RCL. Outputs include structured reasoning chains to enhance interpretability and trust. The contributions of this work are summarized as follows:
@@
    We propose TrioXpert, the first end-to-end framework that unifies AD, FT, and RCL using a collaborative reasoning architecture built on LLMs. The design integrates metrics, logs, and traces through a structured prompting strategy, enabling scalable and explainable multi-task diagnostics.
@@
    We evaluated TrioXpert on two real-world microservice systems and observed consistent improvements over baselines in both performance and interpretability. TrioXpert has also been deployed in Lenovo’s production environment, demonstrating substantial gains in diagnostic efficiency and accuracy. To ensure reproducibility, we release all code, prompts, configurations, and data  [^1] https://anonymous.4open.science/r/TrioXpert-F244.
@@ II-A1 Do Metrics, Logs, and Traces All Carry Diagnostic Value
To understand how various modalities characterize microservice systems, we conducted a qualitative and quantitative analysis of operational data collected from Lenovo’s production environment. Specifically, we collected 20 incident cases and gathered corresponding metrics, raw logs, and traces within a 10-minute window around each incident—reflecting standard incident diagnosis practices in industry deployments [28].
@@ II-C Problem Definition
Typically, OCEs need to collect multimodal data (i.e., metrics, logs, and traces) and perform three core tasks: AD, FT, and RCL. Specifically, the goal of the AD task is to detect whether there are multiple abnormal timestamps within a time window, thereby determining whether the system has encountered an anomaly. The FT task involves selecting the most probable failure type $c_{i}$ from a predefined set of fault categories $\mathbb{C}=\{c_{1},c_{2},...,c_{n}\}$. Finally, the RCL task identifies the most likely root cause $r_{i}$ from all service instances $\mathbb{R}=\{r_{1},r_{2},...,r_{m}\}$.
@@ III-A Overview
In the preprocessing stage, tailored strategies are applied to metrics, logs, and traces to produce structured inputs: a time-series matrix $\mathcal{M}$, a service topology graph $\mathcal{G}$, and filtered logs $\mathcal{L}$ and traces $\mathcal{T}$. Subsequently, three modality-aware pipelines extract numerical and textual features for system status characterization. Finally, an multi-expert collaboration architecture coordinates multiple LLMs to jointly perform downstream incident management tasks, supported by structured reasoning evidence.
Figure 2: The overview of TrioXpert. The framework consists of three modules: (a) Multimodal Data Preprocessing, which filters and prepares metrics, traces, and logs (e.g., aligning by time and extracting relevant subsets); (b) Multi-Dimensional System Status Representation, which derives a numerical feature-based preliminary analysis and textual summaries of logs/traces (via specialized “Log Summarizer” and “Trace Summarizer”); and (c) LLMs Collaborative Reasoning, an multi-expert collaboration architecture comprising a Numerical Expert and Textual Expert that collaborate with an Incident Expert LLM to perform anomaly detection, failure triage, and root cause localization simultaneously.
@@ III-B1 Metrics
Considering the time-series characteristics of metrics, it is organized into a three-dimensional time-series matrix $\mathcal{M}$. The three dimensions of this matrix represent timestamps, service instances (e.g., emailservice-0, logservice-1), and feature channels (e.g., cpu_util, disk_io_rate), respectively. This structured representation not only facilitates subsequent time-series feature extraction but also effectively captures the dynamic changes in system operational status.
@@ III-B2 Logs
In the first stage, the LLM is prompted to extract incident-relevant keywords from the dataset (e.g., generating terms such as “error”, “failure”, or “timeout” in response to the instruction “Identify key terms commonly associated with system failures in the following logs”). This approach eliminates the need for labor-intensive manual curation while ensuring comprehensive coverage of incident-related terminology across different system contexts. The keywords are then employed to filter candidate logs, which are subsequently fed into the second stage for deeper semantic analysis. In the second stage, the LLM executes context-aware filtering through a structured evaluation prompt. This instruction-driven mechanism replaces TF-IDF’s surface-level term frequency analysis with semantic reasoning, enabling the model to: (1) identify causally relevant patterns beyond lexical matching, (2) dynamically weigh diagnostic significance based on contextual evidence, (3) generalize from few examples without handcrafted thresholds [27].
@@
Definition 2: Filtered logs $\mathcal{L}$ are a refined subset of log entries obtained through a two-stage process: keyword-based filtering to retain incident-related logs, followed by LLM prioritization to select distinctive entries.
@@ III-B3 Traces
To identify anomalous invocation chains while preserving end-to-end causality, we implement a type-aware trace filtering mechanism. First, for each invocation type (e.g., HTTP, RPC), we compute a latency threshold as the P95 percentile across all spans of that type, accounting for their distinct latency distributions. Then, when a span’s latency exceeds its type-specific threshold, we recursively include all ancestor spans up to the root node. This ensures two critical properties: (1) all spans contributing to high-latency paths are retained, and (2) no partial invocation chains are preserved, which is essential for accurate root cause localization. Additionally, following the methods of ART [1] and DiagFusion [2], TrioXpert extracts service topology graphs $\mathcal{G}$ from trace data, maintaining the same sampling frequency as the metrics data to describe service instance invocation relationships at any given time.
@@ III-C Multi-Dimensional System Status Representation
To comprehensively capture the state of a microservice system, TrioXpert employs three independent data processing pipelines, targeting both numerical and textual modalities. Each pipeline extracts and encodes salient features from specific modalities, providing structured inputs for downstream reasoning.
@@ III-C1 Numerical Perspective
The numerical pipeline constructs a quantitative system state representation from the time-series matrix $\mathcal{M}$ and service topology graph $\mathcal{G}$. To model heterogeneous dependencies, we adopt the established feature processor architecture from ART [1]: a Transformer Encoder captures channel-wise dependencies, a GRU models temporal dynamics, and GraphSage encodes call-level structural relationships. This processor predicts next-step values of $\mathcal{M}$ and computes deviations from observed metrics. With the deviations above, we generate task-specific preliminary answers using: (a) extreme value theory thresholding for AD, (b) cut-tree-based clustering for FT, and (c) deviation similarity sorting for RCL. All downstream processing follows ART [1] to ensure methodological consistency with standard diagnostic practices. Crucially, the reliability of these preliminary answers stems from self-supervised training, which minimizes the L2 norm of prediction errors to learn robust normal-behavior patterns. Thus, the preliminary answers form a compact numerical representation of the microservice system state.
@@ III-D LLMs Collaborative Reasoning
To address these challenges, TrioXpert adopts the multi-expert collaboration architecture comprising three specialized LLMs: a Numerical Expert, a Textual Expert, and an Incident Expert. Each expert operates on modality-specific inputs and contributes to a collaborative diagnostic process.
@@ III-D2 Coordination Pipeline
TrioXpert adopts the multi-expert collaboration paradigm, leveraging a dispatch-and-aggregation mechanism to fully exploit the specialized strengths of each expert in their respective domains. The pseudo-code for the coordination pipeline is shown in Algorithm 1. Specifically, this mechanism first assigns numerical features $\mathcal{F}_{num}$ and textual features $\mathcal{F}_{txt}$ to the Numerical Expert and Textual Expert, respectively, for independent analysis. Both experts independently generate results and corresponding reasoning evidence for AD, FT, and RCL. Subsequently, the Incident Expert performs a comprehensive analysis of the outputs from the Numerical Expert and Textual Expert, balancing information from different sources to resolve potential conflicts and ultimately producing a unified and reliable diagnostic result $\mathcal{R}_{final}$ along with comprehensive reasoning evidence $\mathcal{E}_{final}$. This dispatch-and-aggregation mechanism not only enables efficient integration of multimodal data but also significantly enhances the accuracy and interpretability of TrioXpert in performing AD, FT, and RCL tasks within complex microservice systems.
@@ 2 $\mathcal{R}_{txt},\mathcal{E}_{txt}\leftarrow$ TextualExpert($\mathcal{F}_{txt}$) ;
Algorithm 1 Coordination Pipeline in TrioXpert
@@ III-D3 Conflict Resolution and Aggregation Policy
Due to potential inconsistencies in incident information across modalities, conflicts may arise between experts. For example, the Numerical Expert may infer a failure type such as “Node CPU overload”, based on time-series metrics, while the Textual Expert may conclude it is a “Container Hardware” issue based on logs and traces. To resolve discrepancies through semantic reasoning rather than numerical fusion, we employ a text-guided weighting mechanism. Specifically, LLM experts are instructed via natural language prompts to prioritize information sources based on their respective reliability.
@@ III-D4 Hallucination Mitigation Design
To address the challenge of hallucination in LLMs, TrioXpert introduces a comprehensive mitigation strategy centered around refined prompt engineering. Specifically, TrioXpert adopts a structured prompt engineering approach grounded in the “RGCIE” principle (i.e., Role, Goal, Constraints, Instructions, Example) to tailor prompts for different LLMs. By explicitly defining roles, goals, instructions, and constraints, along with providing high-quality examples, this approach promotes strict adherence of LLMs to task requirements, thereby significantly reducing the occurrence of hallucinations. As shown in Fig. 4, the prompts are further formalized using YAML format, enabling consistent structuring, easy customization, and programmable parsing across diverse model interfaces. This systematic design not only enhances the precision of model outputs but also aligns them more closely with the intended objectives.
Through this strategy, TrioXpert effectively mitigates the impact of hallucination, contributing to more robust and reliable performance in incident management tasks.
@@ IV Experiments
    RQ3: Is TrioXpert interpretable and practically actionable for real-world incident management?
@@ IV-A1 Datasets
To comprehensively evaluate the performance of TrioXpert in AD, FT, and RCL, we conducted extensive experiments on two datasets, $\mathscr{D}1$ and $\mathscr{D}2$ (shown in Table I). These datasets include three modalities of data (i.e., metrics, logs, traces), and each incident case was annotated by senior industry experts and researchers with extensive experience in the AIOps community. Since these datasets not only encompass multimodal information but also provide high-quality annotations consistent with multi-task labels, they support a comprehensive evaluation of TrioXpert’s performance.
@@ IV-A3 Evaluation Metrics
As described in Section II-C, the goal of TrioXpert is to detect failures, determine failure types, and locate the root cause. To better reflect the performance of the evaluated methods in real-world applications, we designed and selected different evaluation metrics for each task.
AD and FT are both inherently classification tasks, but their objectives differ: AD is a binary classification task used to determine whether a system has experienced a failure, while FT is a multi-class classification task aimed at identifying the type to which the failure belongs. To evaluate the performance of these two tasks, we utilized the following metrics: True Positives (TP), False Positives (FP), and False Negatives (FN). Based on these metrics, we calculated $Precision=\frac{TP}{TP+FP}$ and $Recall=\frac{TP}{TP+FN}$, and subsequently derived the $F1=\frac{2\cdot Precision\cdot Recall}{(Precision+Recall)}$ as a comprehensive evaluation criterion.
For the RCL task, we introduced the $Top@K=\frac{1}{N}\sum_{i=1}^{N}(gt_{i}\in P_{i}[1:K])$ metric to calculate the probability that the true root cause is included within the top $K$ predicted results, where $gt$ represents groundtruth, $P$ stands for predictions and $K$ is the total amount of failures. Additionally, to comprehensively assess model performance, we also employed the $Avg@5=\frac{1}{5}\sum_{i=1}^{5}Top@K$ metric.
@@
IV-A4 Implementations
We implemented TrioXpert using Python 3.10.16 with PyTorch 2.4.0, Transformers 4.46.0, and DGL 2.4.0+cu124. For the backbone LLM, we selected Qwen2.5-7B-Instruct due to its state-of-the-art performance on the Open LLM Leaderboard [31] among models under 10B parameters. Experiments were conducted on a server with 16-core Intel Xeon Gold 5416S CPU, 376GB RAM, and 8 NVIDIA RTX A6000 GPUs (48GB memory each). To ensure result reliability, we repeated each experiment five times and reported the average performance.
IV-B RQ1: Overall Performance